OpenStudy (anonymous):

Integrate the following:

OpenStudy (anonymous):

\[∫ \sin(2x-3)dx\]

OpenStudy (slaaibak):

To make it a bit easier, let's use substitution: \[u = 2x - 3 \rightarrow du = 2 dx \rightarrow dx = {1 \over 2} du\] Now the integral becomes: \[\int\limits {1 \over 2} \sin {u} du = -{1 \over 2}\cos u + C\] substituting u=2x-3 back: \[-{1 \over 2} \cos (2x - 3) + C\]

OpenStudy (anonymous):

i had positive 1/3 not -1/2

OpenStudy (slaaibak):

How did you get a 1/3 ?

OpenStudy (anonymous):

\[-1/3∫ \sin(2-3x)d(2-3x) ... 1/3\cos(2-3x)+c\]

OpenStudy (slaaibak):

but in the problem you stated it's 2x - 3, and not 2 - 3x?

OpenStudy (anonymous):

i meant 2-3x lol sorry!

OpenStudy (hba):

Well if that is the case your answer would be \[\frac{ 1 }{ 3 }\cos(2-3x)+C\] Which you said you got. Therefore,you are correct.

OpenStudy (anonymous):

yay : )